OpenStudy (anonymous):

Solve |x + 2| = 10 i got {-8, 8}?

OpenStudy (anonymous):

One of those is right.

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

Ok, so the trick to doing absolute values is this.. Solve for the absolute value part.

OpenStudy (mathteacher1729):

Here is a graph of this problem. You want to find all values of x such that the blue function is under the red function.

OpenStudy (anonymous):

Then when you take off the absolute value, you get two different equations. \[|a| = b \implies a = b \text{ OR } a=-b\]

OpenStudy (anonymous):

In this case you have \[|x + 2| = 10\] Which means \[x + 2 = 10\]OR \[x+2 = -10\] You have to solve both to find your solutions.